# Searching for GEMS: The Occurrence of Giant Planets orbiting M-dwarfs within 100 pc

### v1

@@ Abstract
We present results from a systematic search for transiting short-period Giant Exoplanets around M-dwarf Stars (GEMS; $P<10$ days, $R_{p}\gtrsim 8~R_{\oplus}$) within a distance-limited 100 pc sample of 149,316 M-dwarfs using TESS-Gaia Light Curve (TGLC) data. This search led to the discovery of one new candidate GEM, following spectroscopic vetting of 12 additional candidates to eliminate astrophysical false positives and refine our occurrence rate estimates. We describe the development and application of the TESS-miner package and associated vetting procedures used in this analysis. To assess detection completeness, we conducted $\sim$ 72 million injection-recovery tests across $\sim$ 26,000 stars with an average of $\sim$3 sectors of data per star, subdivided into early-type (M0–M2.5), mid-type (M2.5–M4), and late-type (M4 or later) M-dwarfs. Our pipeline demonstrates high sensitivity across all M-dwarf subtypes within the injection bounds.
We estimate the occurrence rates of short-period GEMS as a function of stellar mass, and combine our measured occurrence rates with those derived for FGK stars and fit an exponential trend with stellar mass, consistent with core-accretion theory predictions. We find GEMS occurrence rates of $0.067\%\pm 0.047\%$ for early-type M-dwarfs, $0.139\%\pm 0.069\%$ for mid-type, and $0.032\%\pm 0.032\%$ for late-type M-dwarfs, with a mean rate of $0.065^{+0.025}_{-0.027}\%$ across the full M-dwarf sample. We note that while our search spanned $1.0~\mathrm{days}<P<10.0$ days, these occurrence rates were calculated using planets orbiting with $1.0~\mathrm{days}<P<5.0$ days. This work lays the foundation for future occurrence rate investigations for GEMS.
@@
facilities: TESS, HET (HPF), WIYN (NEID), Magellan:Clay (WINERED)software: pickle (Van Rossum, 2020), astropy (Astropy Collaboration et al., 2013, 2018, 2022), barycorpy (Kanodia & Wright, 2018), HxRGproc (Ninan et al., 2018), ipython (Perez & Granger, 2007), K2fov (Mullally et al., 2016), matplotlib (Hunter, 2007), MRExo (Kanodia et al., 2019, 2023a), numpy (Harris et al., 2020), pandas (McKinney, 2010), scipy (Oliphant, 2007; Virtanen et al., 2020), batman (Kreidberg, 2015), lightkurve (Lightkurve Collaboration et al., 2018), wotan (Hippke et al., 2019), WARP (Hamano et al., 2024) , tglc (Han & Brandt, 2023), TRICERATOPS (Giacalone et al., 2021), LEO-vetter (Kunimoto, 2024), juliet (Espinoza et al., 2019), dynesty (Speagle, 2020)
@@ 1 Introduction
M-dwarfs frequently host multiple terrestrial planets (Mulders et al., 2015; Dressing & Charbonneau, 2015; Gaidos et al., 2016; Hardegree-Ullman et al., 2019; Hsu et al., 2020), but the occurrence rate and dominant formation pathway of Giant Exoplanets around M-dwarf Stars (GEMS) remain uncertain (e.g., Endl et al., 2006; Johnson et al., 2010; Maldonado et al., 2019; Schlecker et al., 2022; Gan et al., 2023; Bryant et al., 2023). According to core-accretion theory, giant planets form during the protoplanetary phase through i.) grain condensation, ii.) planetesimal coagulation (Greenberg et al., 1978; Wetherill & Stewart, 1989; Aarseth et al., 1993; Kokubo & Ida, 1996, 2002), iii.) oligarchic growth of planetary embryos (Kokubo & Ida, 1998, 2000), and iv.) runaway gas accretion onto solid cores (Mizuno, 1980; Bodenheimer & Pollack, 1986; Pollack et al., 1996; Ikoma et al., 2000). However, the low disk masses and long Keplerian timescales around M-dwarfs are expected to inhibit this process (Laughlin et al., 2004). Alternatively, GEMS may form via gravitational instability within the protoplanetary disk, whereby regions of the disk implode directly and form gas giants without solid cores (Boss, 1997; Boss & Kanodia, 2023). Regardless of the formation mechanism, GEMS must originate at large orbital separations — due to the implausibility of in-situ formation at their current locations — and subsequently migrate inward to avoid engulfment by the host star (Boss, 2006, 2011). Atmospheric characterization has been proposed as a method to distinguish between formation pathways, but more GEMS discoveries are needed to enable such studies (Helled & Bodenheimer, 2010; Öberg et al., 2011; Madhusudhan et al., 2014; Knierim et al., 2022).
@@
Although transiting GEMS discovered to date have primarily been detected around early M-dwarfs (M0-M2), six GEMS transiting mid M-dwarfs have recently been confirmed: TOI-5205 b (Kanodia et al., 2023b), TOI-3235 b (Hobson et al., 2023), TOI-519 b (Parviainen et al., 2021; Kagetani et al., 2023; Hartman et al., 2023), TOI-4860 b (Almenara et al., 2023; Triaud et al., 2023), TOI-6894 b (Bryant et al., 2025), and TOI-7149 b (Kanodia et al., 2025). Radial velocity (RV) surveys have confirmed the presence of non-transiting GEMS orbiting mid to late M-dwarfs as well (e.g., Morales et al., 2019; Quirrenbach et al., 2022). [^1] We note here that non-transiting GEMS have a $M_{p}\sin{i}$ degeneracy in that their exact orbital inclination cannot be derived from RV observations alone. Therefore, they may still prove to be substellar objects. Furthermore, the Gaia mission has begun to detect candidate GEMS on long-period orbits using the astrometric technique (Holl et al., 2023; Gaia Collaboration et al., 2023a) that can be confirmed using radial velocities (e.g., Stefánsson et al., 2025a). The presence of GEMS orbiting these low-mass stars exceeds the inferred mass budget of solids in protoplanetary disk samples, challenging the predictions of population synthesis models (Burn et al., 2021; Kanodia et al., 2023b).
@@
Before the launch of the Transiting Exoplanet Survey Satellite (TESS; Ricker et al., 2014), the Kepler mission (Borucki et al., 2010) was designed to detect the transits of Earth-sized planets orbiting Sun-like stars, and was only able to achieve the necessary S/N for targets down to $V$-mag $\sim 14$. As a result, the primary target sample included relatively few M-dwarfs, which are generally fainter than this magnitude in the Kepler field of view. Although GEMS-specific studies using ground-based data were able to place upper limits on their occurrence rates (Kovács et al., 2013; Zendejas Dominguez et al., 2013; Obermeier et al., 2016), the discoveries from both Kepler and ground-based surveys were too limited to produce reliable constraints.
@@
Most GEMS detections to date are from RV surveys of nearby M-dwarfs (Endl et al., 2006; Johnson et al., 2010; Bonfils et al., 2013; Maldonado et al., 2019; Sabotta et al., 2021; Schlecker et al., 2022; Pinamonti et al., 2022; Mignon et al., 2025), which have largely confirmed the rarity of these planets by placing upper limits on their occurrence rates. Since 2018, TESS (Ricker et al., 2014) has observed millions of M-dwarfs, uncovering numerous transiting candidate GEMS (Guerrero et al., 2021). Many of these are excellent targets for follow-up observations aimed at confirming their planetary nature and ruling out false positives such as eclipsing binaries (EBs) and brown dwarfs (BDs).
@@
To discover GEMS and measure their occurrence rate, Kanodia et al. (2024) introduced the motivation and framework for the Searching for GEMS survey using TESS data, including the description of a 200 pc M-dwarf sample. Their goal was to enable occurrence rate measurements as a function of stellar mass for a distance-limited sample, which could then be compared to trends in protoplanetary disc mass (Pascucci et al., 2016; Manara et al., 2023) and planetary bulk properties (Müller & Helled, 2024; Kanodia, 2024a). The present study constitutes the first analysis within that framework, focusing on the 100 pc subset of the sample defined by Kanodia et al. (2024). The exact selection criteria and sample boundaries are described in Section 2.
@@
Previous occurrence rate studies of GEMS (Gan et al., 2023; Bryant et al., 2023) have focused on samples which were biased toward earlier-type M-dwarfs due to limiting cuts on TESS sector, stellar mass, and/or stellar effective temperature (see Section 2, Table 1). If core-accretion theory does govern the dominant formation mechanism for GEMS, such a bias would be expected to yield higher occurrence rates (we address this sampling bias in Section 2). These studies were also limited by a high incidence of false positive (FP) candidates, often arising from EBs. High FP rates also contribute to the overestimation of the occurrence rates. In this work, we derive occurrence rates for GEMS across the full M-dwarf mass range and perform astrophysical false positive validation via RV follow-up for all but two of our candidates (see Section 4.6) to reduce the number of false positives and ensure a robust occurrence rate analysis.
@@ 2 Sample Characteristics
Figure 1: The stellar characteristics of the GEMS 100 pc sample compared with those of Bryant et al. (2023) in orange and Gan et al. (2023) in green. They are as follows: (a) stellar mass, (b) effective temperature, and the cumulative distribution functions (CDFs) of (c) distance, (d) $G$ mag, (e) $K_{S}$ mag, and (f) $T$ mag. In the latter three, the GEMS 100 pc sample (in dark blue) is divided into early (solid), mid (dashed), and late (dotted).
Table 1: A comparison of (a) the cuts used to define, and (b) the descriptive metrics of Bryant et al. (2023), Gan et al. (2023), and this work.
@@
Bryant et al. (2023) Gan et al. (2023) This work
@@
aafootnotetext: https://archive.stsci.edu/hlsp/tess-spoc (TESS-SPOC; Caldwell et al. 2020)
bbfootnotetext: https://archive.stsci.edu/hlsp/qlp (QLP; Huang et al. 2020)
ccfootnotetext: https://archive.stsci.edu/hlsp/tglc (TGLC; Han & Brandt 2023a)
We draw our M-dwarf sample from the 200 pc distance-limited catalog described in Kanodia et al. (2024), and observed in the TESS primary mission (PM; TESS sectors 1 – 26) and first extended mission (EM1; TESS sectors 27 – 55). The 100 pc subset is designated as a proving ground for the methods in this work, with the full 200 pc sample reserved for a future search. We queried our targets from Gaia DR3 (Gaia Collaboration et al., 2023b), according to the same parallax error, quality, color, and $M_{K_{s}}$ criteria [^2] We did not impose a cut on Gaia’s Renormalized Unit Weight Error (RUWE) or any other astrometric quality metric, as doing so could have biased our faint sample toward brighter, better-fitted targets. as laid out in Section 4.3.1 of Kanodia et al. (2024). For the 100 pc sample, we include M-dwarfs with:
@@
    Null flags indicating that the target is not in the Gaia DR3 qso_candidates, galaxy_candidates, or nss_two_body_orbit tables (Gaia Collaboration et al., 2023a).
These cuts resulted in a total of 149,316 unique stars. We used the absolute $M_{K_{S}}$ bounds from Kanodia et al. (2024) to split this sample into $27,995$ early (<< M2.5; $4.7<M_{K_{S}}<6.0$), $37,948$ mid (M2.5 – M4; $6.0<M_{K_{S}}<7.1$), and $79,669$ late M-dwarfs (>> M4; $7.1<M_{K_{S}}<10.1$). The masses for all stars in our sample were calculated using the $M_{K}$ relations from Equation 4 in Mann et al. (2019), and the radii and effective temperatures ($T_{\text{eff}}$) using Equation 4 in Mann et al. (2015) (see Mann et al. (2015) Table 1 for the coefficients associated with the calculation of $R_{*}$ and $T_{\text{eff}}$). [^3] Our sample slightly exceeds the calibrated range of the $M_{K_{S}}$$T_{\text{eff}}$ relation. For stars with $9.1<M_{K_{S}}<10.1$, we assign $T_{\text{eff}}=2700~\mathrm{K}$ by convention, which produces the spike seen in Figure 1b. We do not correct for reddening when creating the sample. While extinction is generally minimal within 100 pc (Reis et al., 2011; Gontcharov, 2017), we note that some regions may still be affected by modest reddening, which could introduce small biases in $M_{K}$-based stellar parameters. A table of our 100 pc sample, along with a table of all targets surviving until manual vetting (see Section 4.2), is available online444 https://zenodo.org/records/15738847.
Figure 1 compares our sample to those from Gan et al. (2023) and Bryant et al. (2023), two previous TESS-based searches for transiting GEMS. Table 1 summarizes key differences in sample selection, cadence, and sector coverage. Our sample most closely resembles that of Bryant et al. (2023). Gan et al. (2023) notably extended their search to 200 pc, beyond the 100 pc limit adopted here and by Bryant et al. (2023). As shown in Figure 1a and b, our sample includes a significantly more complete set of low-mass stars ($M_{\star}/M_{\odot}<0.4$) than either prior study. This is largely due to the fact that the MIT Quick Look Pipeline used by Gan et al. (2023) (QLP; Huang et al. (2020)) and the TESS Science Processing Operations Center Pipeline used by Bryant et al. (2023) (TESS-SPOC; Caldwell et al. (2020)) have more restrictive magnitude limits than the TESS-Gaia Light Curve (TGLC; Han & Brandt, 2023a) pipeline which we use (see Section 3.1.)
@@
We applied the same selection criteria that define our target sample to the NASA Exoplanet Archive [^5] https://exoplanetarchive.ipac.caltech.edu/ (Christiansen et al., 2025) to retrieve all confirmed GEMS, adopting $0.7<R_{p}/R_{J}<1.5$ and $1.0<P/\mathrm{day}<10.0$ as our working definition. [^6] The lower radius limit corresponds to $R_{p}-\sigma(R_{p})\geq 8.0~R_{\oplus}$; we quote radii in $R_{J}$ to one decimal place hereafter. As of this writing, the Archive lists 20 confirmed GEMS that (i) lie within 200 pc, (ii) satisfy our magnitude cuts, and (iii) were observed in PM and/or EM1. We also included TOI-6894 b ($d\simeq 73$ pc), recently confirmed by Bryant et al. (2025). We selected the thresholds within our pipeline (Section 3, Section 4) so that all of the confirmed GEMS advanced to the manual vetting stage.

### v3

@@ Abstract
We present results from a systematic search for transiting short-period Giant Exoplanets around M-dwarf Stars (GEMS; $P<10$ days, $R_{p}\gtrsim 8~R_{\oplus}$) within a distance-limited 100 pc sample of $149,316$ M-dwarfs using TESS-Gaia Light Curve (TGLC) data. We describe the development and application of the TESS-miner package and associated vetting procedures used in this analysis. To assess detection completeness, we conducted $\sim$72 million injection-recovery tests across $\sim$26,000 stars with an average of $\sim$3 sectors of data per star, subdivided into early-type (M0–M2.5), mid-type (M2.5–M4), and late-type (M4 or later) M-dwarfs. Our pipeline demonstrates high sensitivity across all subtypes within the injection bounds.
We estimate the occurrence rates of short-period GEMS as a function of stellar mass, and combine our measured rates with those derived for FGK stars, fitting an exponential trend with stellar mass, consistent with core-accretion theory predictions. We find GEMS occurrence rates of $0.118\%\pm 0.068\%$ for early-type M-dwarfs, $0.153\%\pm 0.069\%$ for mid-type, and $0.036\%\pm 0.024\%$ for late-type M-dwarfs, with a mean rate of $0.068\%\pm 0.024\%$ across the full sample. While our search spanned $1.0~\mathrm{days}<P<10.0$ days, these rates were calculated using planets orbiting with $1.0~\mathrm{days}<P<5.0$ days. This work establishes the basis for future occurrence rate studies of transiting GEMS.
@@
facilities: TESS, HET (HPF), WIYN (NEID), Magellan:Clay (WINERED)software: pickle (van1995python), astropy (astropy_2013; astropy_collaboration_astropy_2018; astropy_2022), barycorpy (kanodia_python_2018), HxRGproc (ninan_habitable-zone_2018), ipython (perez_ipython_2007), K2fov (2016ascl.soft01009M), matplotlib (hunter_matplotlib:_2007), MRExo (kanodia_mass-radius_2019; kanodia_beyond_2023), numpy (harris_array_2020), pandas (mckinney-proc-scipy-2010), scipy (oliphant_python_2007; scipy), batman (kreidberg_batman_2015), lightkurve (lightkurve), wotan (Wotan), WARP (2024PASP..136a4504H) , tglc (han_tess-gaia_2023), TRICERATOPS (giacalone_vetting_2021), LEO-vetter (Kunimoto_2024), juliet (Espinoza2019), dynesty (speagle_dynesty_2020)
@@ 1 Introduction
M-dwarfs frequently host multiple terrestrial planets (Mulders2015; dressing_occurrence_2015; gaidos_they_2016; hardegree_ullman_kepler_2019; hsu_occurrence_2020), but the occurrence rate and dominant formation pathway of Giant Exoplanets around M-dwarf Stars (GEMS) remain uncertain (e.g., endl_exploring_2006; johnson_giant_2010; maldonado_connecting_2019; schlecker_rv-detected_2022; gan_occurrence_2023; bryant_occurrence_2023). According to core-accretion theory, giant planets form during the protoplanetary phase through i.) grain condensation, ii.) planetesimal coagulation (Greenberg1978_collisional_evol; Wetherill1989; Aarseth1993; Kokubo1996; Kokubo02), iii.) oligarchic growth of planetary embryos (Kokubo1998; Kokubo2000), and iv.) runaway gas accretion onto solid cores (Mizuno1980; bodenheimer_calculations_1986; Pollack96; ikoma_formation_2000). However, the low disk masses and long Keplerian timescales around M-dwarfs are expected to inhibit this process (laughlin_core_2004). Alternatively, GEMS may form via gravitational instability within the protoplanetary disk, whereby regions of the disk implode directly and form gas giants without solid cores (boss_giant_1997; boss_forming_2023). Regardless of the formation mechanism, GEMS must originate at large orbital separations — due to the implausibility of in-situ formation at their current locations — and subsequently migrate inward to avoid engulfment by the host star (boss_formation_2006; boss_formation_2011). Atmospheric characterization has been proposed as a method to distinguish between formation pathways, but more GEMS discoveries are needed to enable such studies (helled_metallicity_2010; oberg_effects_2011; madhusudhan_toward_2014; knierim_constraining_22).
@@
Although transiting GEMS discovered to date have primarily been detected around early M-dwarfs (M0-M2), six GEMS transiting mid M-dwarfs have recently been confirmed: TOI-5205 b (kanodia_toi-5205b_2023), TOI-3235 b (hobson_toi-3235_2023), TOI-519 b (parviainen_toi-519_2021; kagetani_mass_2023; hartman_toi_2023), TOI-4860 b (almenara_toi-4860_2023; triaud_m_2023), TOI-6894 b (Bryant2025), and TOI-7149 b (kanodia_toi-7149_2025). Radial velocity (RV) surveys have confirmed the presence of non-transiting GEMS orbiting mid to late M-dwarfs as well (e.g., morales_giant_2019; quirrenbach_carmenes_2022-1). [^1] We note here that non-transiting GEMS have a $M_{p}\sin{i}$ degeneracy in that their exact orbital inclination cannot be derived from RV observations alone. Therefore, they may still prove to be substellar objects. Furthermore, the Gaia mission has begun to detect candidate GEMS on long-period orbits using the astrometric technique (holl_gaia_2023; gaia_collab_2023_stellar_mult) that can be confirmed using radial velocities (e.g., Stefansson2025). The presence of GEMS orbiting these low-mass stars exceeds the inferred mass budget of solids in protoplanetary disk samples, challenging the predictions of population synthesis models (burn_new_2021; kanodia_toi-5205b_2023).
@@
Before the launch of the Transiting Exoplanet Survey Satellite (TESS; Ricker14), the Kepler mission (Borucki10) was designed to detect the transits of Earth-sized planets orbiting Sun-like stars, and was only able to achieve the necessary S/N for targets down to $V$-mag $\sim 14$. As a result, the primary target sample included relatively few M-dwarfs, which are generally fainter than this magnitude in the Kepler field of view. Although GEMS-specific studies using ground-based data were able to place upper limits on their occurrence rates (kovacs_hot_2013; Zendejas2013; Obermeier2016), the discoveries from both Kepler and ground-based surveys were too limited to produce reliable constraints.
@@
Most GEMS detections to date are from RV surveys of nearby M-dwarfs (endl_exploring_2006; johnson_giant_2010; bonfils_harps_2013; maldonado_connecting_2019; sabotta_carmenes_2021; schlecker_rv-detected_2022; pinamonti_hades_2022; mignon2025), which have largely confirmed the rarity of these planets by placing upper limits on their occurrence rates. Since 2018, TESS (Ricker14) has observed millions of M-dwarfs, uncovering numerous transiting candidate GEMS (guerrero_tess_2021). Many of these are excellent targets for follow-up observations aimed at confirming their planetary nature and ruling out false positives such as eclipsing binaries (EBs) and brown dwarfs (BDs).
@@
To discover GEMS and measure their occurrence rate, kanodia2024 introduced the motivation and framework for the Searching for GEMS survey using TESS data, including the description of a 200 pc M-dwarf sample. Their goal was to enable occurrence rate measurements as a function of stellar mass for a distance-limited sample, which could then be compared to trends in protoplanetary disc mass (pascucci_steeper_2016; Manara23) and planetary bulk properties (Muller_2024; Kanodia_2025). The present study constitutes the first analysis within that framework, focusing on the 100 pc subset of the sample defined by kanodia2024. The exact selection criteria and sample boundaries are described in Section 2.
@@
Previous occurrence rate studies of GEMS (gan_occurrence_2023; bryant_occurrence_2023) have focused on samples which were biased toward earlier-type M-dwarfs due to limiting cuts on TESS sector, stellar mass, and/or stellar effective temperature (see Section 2, Table 1). If core-accretion theory does govern the dominant formation mechanism for GEMS, such a bias would be expected to yield higher occurrence rates (we address this sampling bias in Section 2). These studies were also limited by a high incidence of false positive (FP) candidates, often arising from EBs. High FP rates also contribute to the overestimation of the occurrence rates. In this work, we derive occurrence rates for GEMS across the full M-dwarf mass range and perform astrophysical false positive validation via RV follow-up for all of our candidates (see Section 4.6) to remove false positives and ensure a robust occurrence rate analysis.
@@ 2 Sample Characteristics
Figure 1: The stellar characteristics of the GEMS 100 pc sample compared with those of bryant_occurrence_2023 in orange and gan_occurrence_2023 in green. They are as follows: (a) stellar mass, (b) effective temperature, and the cumulative distribution functions (CDFs) of (c) distance, (d) $G$ mag, (e) $K_{S}$ mag, and (f) $T$ mag. In the latter three, the GEMS 100 pc sample (in dark blue) is divided into early (solid), mid (dashed), and late (dotted).
Table 1: A comparison of (a) the cuts used to define, and (b) the descriptive metrics of bryant_occurrence_2023, gan_occurrence_2023, and this work.aafootnotetext: https://archive.stsci.edu/hlsp/tess-spoc (TESS-SPOC; caldwell_tess_2020)bbfootnotetext: https://archive.stsci.edu/hlsp/qlp (QLP; huang_photometry_2020)ccfootnotetext: https://archive.stsci.edu/hlsp/tglc (TGLC; Han2023)
@@
bryant_occurrence_2023 gan_occurrence_2023 This work
@@
We draw our M-dwarf sample from the 200 pc distance-limited catalog described in kanodia2024, and observed in the TESS primary mission (PM; TESS sectors 1 – 26) and first extended mission (EM1; TESS sectors 27 – 55). The 100 pc subset is designated as a proving ground for the methods in this work, with the full 200 pc sample reserved for a future search. We queried our targets from Gaia DR3 (gaia_collaboration_gaia_2023), according to the same parallax error, quality, color, and $M_{K_{s}}$ criteria [^2] We did not impose a cut on Gaia’s Renormalized Unit Weight Error (RUWE) or any other astrometric quality metric, as doing so could have biased our faint sample toward brighter, better-fitted targets. as laid out in Section 4.3.1 of kanodia2024. Following Moe2021, we avoid a simple magnitude-limited search, which can bias a sample toward close, bright binaries. Such objects are already removed from RV target lists, artificially boosting hot-Jupiter occurrence rates relative to transit-based surveys such as this one. For the 100 pc sample, we include M-dwarfs with:
@@
    Null flags indicating that the target is not in the Gaia DR3 qso_candidates, galaxy_candidates, or nss_two_body_orbit tables (gaia_collab_2023_stellar_mult).
These cuts resulted in a total of 149,316 unique stars. We used the absolute $M_{K_{S}}$ bounds from kanodia2024 to split this sample into $27,995$ early (<< M2.5; $4.7<M_{K_{S}}<6.0$), $37,948$ mid (M2.5 – M4; $6.0<M_{K_{S}}<7.1$), and $79,669$ late M-dwarfs (>> M4; $7.1<M_{K_{S}}<10.1$). The masses for all stars in our sample were calculated using the $M_{K}$ relations from Equation 4 in Mann19, and the radii and effective temperatures ($T_{\text{eff}}$) using Equation 4 in mann_how_2015 (see mann_how_2015 Table 1 for the coefficients associated with the calculation of $R_{*}$ and $T_{\text{eff}}$). [^3] Our sample slightly exceeds the calibrated range of the $M_{K_{S}}$$T_{\text{eff}}$ relation. For stars with $9.1<M_{K_{S}}<10.1$, we assign $T_{\text{eff}}=2700~\mathrm{K}$ by convention, which produces the spike seen in Figure 1b. We do not correct for reddening when creating the sample. While extinction is generally minimal within 100 pc (Reis2011; Gontcharov2017), we note that some regions may still be affected by modest reddening, which could introduce small biases in $M_{K}$-based stellar parameters. A table of our 100 pc sample, along with a table of all targets surviving until manual vetting (see Section 4.2), is available online [^4] https://zenodo.org/records/15738847.
Figure 1 compares our sample to those from gan_occurrence_2023 and bryant_occurrence_2023, two previous TESS-based searches for transiting GEMS. Table 1 summarizes key differences in sample selection, cadence, and sector coverage. Our sample most closely resembles that of bryant_occurrence_2023. gan_occurrence_2023 notably extended their search to 200 pc, beyond the 100 pc limit adopted here and by bryant_occurrence_2023. As shown in Figure 1a and b, our sample includes a significantly more complete set of low-mass stars ($M_{\star}/M_{\odot}<0.4$) than either prior study. This is largely due to the fact that the MIT Quick Look Pipeline used by gan_occurrence_2023 (QLP; huang_photometry_2020) and the TESS Science Processing Operations Center Pipeline used by bryant_occurrence_2023 (TESS-SPOC; caldwell_tess_2020) have more restrictive magnitude limits than the TESS-Gaia Light Curve (TGLC; Han2023) pipeline which we use (see Section 3.1.)
@@
We applied the same selection criteria that define our target sample to the NASA Exoplanet Archive [^5] https://exoplanetarchive.ipac.caltech.edu/ (Christiansen2025) to retrieve all confirmed GEMS, adopting $0.7<R_{p}/R_{J}<1.5$ and $1.0<P/\mathrm{day}<10.0$ as our working definition. [^6] The lower radius limit corresponds to $R_{p}-\sigma(R_{p})\geq 8.0~R_{\oplus}$; we quote radii in $R_{J}$ to one decimal place hereafter. As of this writing, the Archive lists 20 confirmed GEMS that (i) lie within 200 pc, (ii) satisfy our magnitude cuts, and (iii) were observed in PM and/or EM1. We also included TOI-6894 b ($d\simeq 73$ pc), recently confirmed by Bryant2025. We selected the thresholds within our pipeline (Section 3, Section 4) so that all of the confirmed GEMS advanced to the manual vetting stage.